A 3.60 g bullet moves with a speed of 153 m/s perpendicular to the Earth's magnetic field of 4.80
Answer #1

The force on the bullet will be the Lorentz force:

F = q*(V x B)

where V = velocity and B the magnetic field. x is the vector cross product.

Since you are not concerned with the direction of the deflection and V and B are perpendicular, this simplifies to

F = q*V*B

This will cause the bullet to accelerate at right angles to its velocity at the rate a = F/m:

a = q*V*B/m

After a time t, the deflection will be

s = 0.5*a*t